Sum of n terms of the sequence [ {1/1+1^2+1^4}+{2/1+2^2+2^4}+{3/1+3^2+3^4}+-----------------------

Grade:

1 Answers

Swapnil Saxena
12 years ago

Σ ( n + n2 + n4) = Σn + Σ n2 + Σ n4 

Σ n = n(n+1)/2

Σ n2 = n(n+1)(2n+1)/6

Σ n4 = n(n+1)(2n+1)(3n2+3n-1)/30

So sum = {n(n+1)/2} + {n(n+1)(2n+1)/6} + {n(n+1)(2n+1)(3n2+3n-1)/30}

So the sum can be calculated in this way.